# Topological Phases, Criticality, and Mixed State Order in a Hubbard Quantum Simulator

### v1

@@ III Coupling an even number of chains removes the topological transition
which results in a gapped path between the stacks of two CSPT-$1$ and CSPT-$2$ phases (Fig. 3a top). Here, $a_{x,\mathsf{T}/\mathsf{B}}^{(\dagger)}$ labels the bosonic annihilation (creation) operator in the top/bottom layer. Crucially, the above term preserves the protecting $\mathbb{Z}_{2}$ parity symmetry generated by Eq. (2), as well as site-centered inversion.
To perform a quantum simulation of the above coupling in a number-conserving system, we map the unity-filling CSPTs to the half-filling ones (Methods) and perform a particle-hole transformation on the bottom layer, treating the physical bosons present in the bottom layer as holes (Fig. 3a bottom). In this representation, the pairing term ($\Delta_{\perp}$) becomes an inter-layer hopping term ($t_{\perp}$), which can be implemented experimentally. The staggered potential is created by a diagonal superlattice, resulting in a checkerboard pattern.
@@
We start deep in the CSPT-1 states of both chains with a negligible $t_{\perp}$. Then, to probe the gap closing in the phase transition, we ramp across the critical point of a CSPT chain, $\mu_{S}=0$. We ramp into a regime where the ground state of each decoupled chain is the CSPT-2 state before imaging. The numerical contour plot in Fig. 3b shows that, with increasing $t_{\perp}$ between two chains, the many-body energy gap between the ground and the first excited states increases at $\mu_{S}=0$, removing the quantum criticality. The ramps in $\mu_{S}$ and $t_{\perp}$ follow the differently shaded red lines at a constant rate for the $\mu_{S}$ ramp.
@@
At low interchain tunneling $t_{\perp}\rightarrow 0$, the ramp cannot be adiabatic as we cross the vanishing energy gap at the quantum critical point. The system thus heats, resulting in a low fidelity for the final state with respect to the CSPT-2 ground state (Fig. 3e left). Consistently, the CSPT-1 order parameter does not vanish in this regime. In contrast, at larger values of $t_{\perp}$, even though the total path traveled in the phase diagram is longer—because we have to change $t_{\perp}$ while simultaneously changing $\mu_{S}$ for the same total sweep time—we observe a more faithful preparation of CSPT-2 state. The larger energy gap allows us to ramp across the phase diagram with less heating, which is also reflected in the final temperature of the prepared CSPT-2 states (Fig. 3c). The energy of the final state can be well approximated by the energy of the classical state because, at the readout stage, the intrachain tunneling ($t$) is much smaller than other energy scales ($\mu_{S}$, $U$, and $V$). Based on each snapshot, we compute the energy of the final state and obtain a histogram of the state’s distribution. We then fit a temperature and observe the shift towards lower temperature, as $t_{\perp}$ is increased.
@@ IV Mixed State Order from Average Symmetry
Figure 4: Average symmetry protects CSPTs subject to local disorder. a, To demonstrate the average symmetry property of the CSPTs, we intentionally add different programmable chemical potential disorder, changing the original light gray lattice to the purple lattice. In the lower panel, the order parameters of a single chemical potential disorder realization (bootstrapping standard error) no longer show clear and distinct phases. The left side lower panel shows the order profile in space for different staggered potentials, while the right side lower panel integrates over the space to plot string order parameters. b, The histograms of the distribution of the $p_{1}$ and $p_{2}$ are plotted for the different disorder realizations. The disorder causes strong deformation in the order parameters for each disorder realization. c, Since the CSPTs are protected by average symmetry, once we average over the different disorder realizations, we obtain a clean signal of two distinct phases (standard error of the mean). The order parameters do not reach exactly zero due to finite imaging fidelity, finite size effects, and the degree of approximation of the particle-hole symmetry (Methods), but still can distinguish between the two phases.
@@
Thus far, we have demonstrated the presence of two distinct CSPTs in our experiment and exhibited a hallmark feature of the transition between them, i.e., the disappearance of the transition upon stacking. Crucially, our results depended on the presence of both protecting symmetries of the CSPT: the crystalline site-centered inversion symmetry and the boson parity symmetry. To further substantiate our claim that crystalline site-centered inversion symmetry is essential, we explicitly break this symmetry by implementing a programmable chemical potential disorder pattern shown at the top of Fig. 4a, whose magnitude is comparable to the largest term in the Hamiltonian (peak-to-peak $\approx U$, see Methods). In our experiment, we adiabatically prepare the ground state of the disordered staggered dipolar Bose-Hubbard model $t\ll U,\mu_{S}$ by ramping down the tunneling from a superfluid while ramping up the chemical potential disorder pattern. For any given realization of the disorder, we repeat the experiment thousands of times (Methods) to sample the wavefunction. This yields the parity string order parameters $p_{1}$ and $p_{2}$ shown in Fig. 4a bottom, which show no signature of a transition with varying $\mu_{S}$, indicating that the transition is removed when inversion symmetry is broken.
However, let us now consider averaging over disorder realizations. We draw disorder realizations $\mathcal{D}$ from a distribution $p_{\mathcal{D}}$ such that disorder patterns related by site-centered inversion are equally likely. Then the ground state for each disorder realization |$\ket{\psi_{\mathcal{D}}}$ is probed by the following
@@
Crucially, probing the above is equivalent to experimentally probing the expected value of the parity string operator in the mixed state $\rho=\sum_{\mathcal{D}}p_{\mathcal{D}}\ket{\psi_{\mathcal{D}}}\bra{\psi_{\mathcal{D}}}$ | with $\mathbb{E}[\mathcal{P}_{x_{0}}(r)]=\text{Tr}\left(\rho\,\hat{\mathcal{P}}_{x_{0}}(r)\right)$. Although for each realization there is no sharp distinction between regimes $\mu_{S}\ll U$ and $\mu_{S}\gg U$ (as seen by the spread in parity values present in the histograms in Fig. 4b), upon averaging over disorder, a sharp distinction is restored (Fig. 4c). This is because the crystalline SPTs we study are also average SPTs, i.e., topological phases of the mixed states that arise due to the randomness of the disorder [14, 15, 16, 17, 18, 19]. Such average SPTs can be shown in theory to be protected by the $\mathbb{Z}_{2}$ parity symmetry along with an average site-centered inversion symmetry of the mixed state corresponding to ^$\hat{\mathcal{I}}\rho\hat{\mathcal{I}}^{\dagger}=\rho$, where $\hat{\mathcal{I}}$ is the unitary operator implementing the action of inversion. Our observations suggest that symmetry protection in realistic quantum systems can persist beyond pure-state assumptions, opening avenues for robust topology under decoherence, finite entropy, or driven-dissipative conditions.
V The Mott-to-Haldane insulator Transition
We conclude by providing evidence for the existence of the Haldane SPT appearing within our system and further elucidate its connection to the crystalline SPTs explored in this work. To do so, let us note that, in the absence of the staggered potential ($\mu_{\text{S}}=0$), the phase diagram of the dipolar Bose-Hubbard model is theoretically predicted to contain at least three insulating phases as a function of the on-site interaction strength $U$ and the characteristic strength of the long-range interactions $V$ (schematically reproduced in Fig. 5a) [21, 62]. Two of these are the Mott insulator (MI) and the charge density wave (CDW), the latter of which spontaneously breaks the translation symmetry $\mathbb{Z}$ of the model at $\mu_{\text{S}}=0$ down to $2\mathbb{Z}$. The last phase is the Haldane Insulator (HI)—an entangled SPT phase that is separated from the Mott insulator by a quantum phase transition in the presence of the crystalline bond-centered inversion symmetry and parity symmetry of the dipolar Bose-Hubbard model at $\mu_{\text{S}}=0$ [32]. In this sense, the HI is a CSPT phase with protected quantum entanglement.
Naturally, CSPT-1 and MI correspond to the same boson pattern and hence the same phase of matter. More intriguingly, the CSPT-2 phase we encountered for $\mu_{S}\neq 0$ adiabatically connects to HI as we take $\mu_{S}\to 0$ [6]. To gain an intuitive understanding of this relationship, we use the “dressed quasi-particle” picture of the CSPTs introduced in Section II. In particular, HI can be thought of as a case where bosons in our system are each delocalized across the bonds of the lattice as shown in Fig. 5a—a configuration that is symmetric under the parity, site- and bond-centered inversion symmetries of the model. Concretely, a model wavefunction for this phase is given by,
@@ V The Mott-to-Haldane insulator Transition
which captures the bond delocalization. If we turn on the staggering $\mu_{S}\neq 0$, these delocalized bosons gradually move together to form the CSPT-2 state as schematically shown in Fig. 1a. Moreover, the direct topological phase transition between CSPT-1 and CSPT-2 (for $\mu_{S}\neq 0$) is in the same universality class as the transition between MI and HI (for $\mu_{S}=0$) [61]. We can thus explore the same topological phase transition in two very different energetic regimes.
@@
Observe that this is similar to the parity string operator we used for CSPT-1 and CSPT-2, however, now there an endpoint dressing $\delta\hat{n}=\hat{n}-1$. The reason for this is that without the dressing, the expected parity would vanish since the bond-delocalized bosons are equally likely to be inside as outside the region of interest. The dressing accounts for this, and the nonzero value of $\mathcal{S}_{x_{0}}(r)$ is predicted to be positive (for odd lengths), consistent with the CSPT-2 phenomenology. Note that this endpoint operator $\delta n$ transforms to $-\delta n$ under the approximate charge-conjugation symmetry of the chain—it is precisely the transformation law of this endpoint decoration that serves as a topological invariant of the HI [10, 11, 12].
To summarize, the hallmark signature of the Haldane insulator is then a contrast between the long-range order of $\mathcal{S}_{x_{0}}(r)$ along with the parity string operator $\mathcal{P}_{x_{0}}(r)$ decaying to zero at long distances. To study the transition between the HI and MI, we turn off the staggered potential and adiabatically load our atoms into the lattice while setting magnetic fields to probe different on-site interactions $U$ [65]. Since doublon-hole fluctuation fractions limit the magnitude of Haldane string order parameter $\mathcal{S}_{x_{0}}(r)$ and it oscillates as a function or $r$, we normalize it as $\widetilde{\mathcal{S}}=(-1)^{r}\eta\mathcal{S}$ with $\eta=1/\langle|\delta n_{i}||\delta n_{i+r}|\rangle$ [42] computed for the state with the maximum string order parameter in the probed parameter space ($V/t\approx 3$). The result is shown in Fig. 5b and reveals two regimes of the phase diagram, one with Haldane-like correlations and the other with Mott-like correlations. In our finite-size system (10 sites), the order parameters are not expected to exhibit sharp features. Nevertheless, since the order parameters we observe are consistent with the simulated results for our finite-size system, Fig. 5b provides evidence of a tunable, interacting Haldane insulator-like SPT state in an analog quantum simulator.
Figure 5: Symmetry protected topological Haldane insulator. A well-known SPT phase in the dipolar Bose-Hubbard model is the Haldane Insulator (HI), which can be connected to the CSPT-2 phase. a, Once the staggered field is turned off, the CSPT-1 and -2 phases become Mott and Haldane insulators in the presence of off-site interactions ($V$). A schematic wavefunction for the HI described in Eq. (6) is shown in the inset. b, We tune the on-site interaction $U$ without the staggered field $\mu_{S}$ at fixed $V/t\approx 3$. The order parameters show good agreement with simulated results. The vertical error bars are obtained from bootstrapping standard error; the horizontal error bars are estimated standard error.
@@ VI Conclusion and outlook
Our study reveals that the topological criticality at the transition between the Haldane insulator (HI) and the Mott insulator (MI) can more easily be accessed between crystalline symmetry-protected topological (CSPT) phases. This underscores the power of tunable quantum simulators, which grant access to a broad parameter space and enable the study of quantum criticality in otherwise inaccessible regimes. While local observables fail to distinguish these phases away from the atomic limit, we identified nonlocal string order parameters for the HI-MI and CSPT transitions. Both transitions are predicted to exhibit stacking behavior, which defines the group structure of SPTs and affirms their invertible nature. We demonstrated this property through dynamic ramping across the CSPT critical point. In the CSPT setting, we also observed the emergence of mixed state order, which generalizes topological invariants into mixed states.
Our uncovering of quantum phase transitions between 1D phases admitting atomic limits raises the possibility of analogous phenomena in two or higher dimensions—an open direction for both theoretical and experimental exploration. These transitions may help chart adiabatic pathways between topologically trivial and non-trivial phases, motivating novel approaches to the discovery of phases of matter [66] starting from experimentally accessible product states such as paramagnets [67, 68]. In parallel, our demonstration of mixed-state order raises compelling questions about its stability, the nature of intervening phases, and how mixed-state perspectives might refine our understanding of real-world quantum systems. Concretely, this same approach could be used to realize topological phases with average crystalline symmetries which can only exist in the presence of disorder [19, 69]. Moreover, it would be interesting for future work to compare the state preparation fidelity of SPT phases via explicit symmetry breaking versus the stacking property demonstrated in our work. The latter might be especially useful for preparing two-dimensional SPT phases protected by, e.g., particle parity symmetry [70], which would be difficult to break explicitly. Our results emphasize that even seemingly simple models can harbor nontrivial topology, reinforcing the idea that topological tools are not just abstract concepts but practical frameworks for identifying and organizing quantum phases.
@@ Methods
Figure 6: Chemical potential. a, Our microscope allows three types of potential projections: we project superlattices using the blue accordion lattice, project arbitrary potentials using DMD, and project relatively sharp repulsive walls with blue-detuned elliptical beams. b, A schematic chemical potential landscape along the 1D chain for a 10-site system. With DMD to compensate for harmonic confinement and repulsive walls, the central 10 sites (shaded in purple) give us a well-defined system. The rising wall marked in purple dots follows the chemical potential listed in the text. c, To increase the data acquisition rate, we parallel simulate multiple copies of chains. In the top subfigure, we raise the lattice deep to reduce interchain tunneling and tilt the magnetic field to minimize interchain dipole-dipole interactions. In the bottom subfigure, we use an additional superlattice (light blue) to create isolated pairs of chains.
@@
Figure System size (sites) Atom number required Atom number DS Photon count DS Spin-half mapping DS Overall DS Smallest sample size Average sample size Total sample size
@@
In the main text, we presented that decoupled chains have two topologically distinct CSPTs, but two tunnel-coupled chains trivialize the topology. Using ED with three by eight sites and four by six sites, we further numerically simulate higher numbers of coupled chains to demonstrate the $\mathbb{Z}_{2}$ properties as shown in Fig. 10.
@@
Strictly speaking, the system only has a site-centered inversion symmetry for odd-length systems. Nevertheless, for the bulk phase of matter properties, this distinction between even and odd system sizes should not play a role.
@@
While local observables such as the average site occupation respond to the staggered chemical potential, they cannot serve as order parameters for distinguishing the two crystalline SPTs. This is because the Hamiltonian already breaks translational symmetry explicitly, so density modulations are always present—even in trivial phases—and they vary smoothly across the transition. To serve as an order parameter, the observable has to be zero outside the phase and non-zero inside the phase. In contrast, the topological invariant distinguishing the phases (e.g., the sign of the inversion-symmetric parity string) changes singularly and is protected by symmetry and a finite gap [6, 10].
@@
In the main text, we claimed that the crystalline SPT phases we study persist provided that the site-centered inversion symmetry of the chain is preserved on average. Here, we provide a theoretical proof for this claim. Before delving into the proof, let us review some of the foundational concepts in the many-body physics of mixed states that will be necessary for our analysis.
@@
A central property of pure state SPTs is that they are short-range entangled (SRE)—i.e.  they can be prepared from a product state by finite time evolution with a local Hamiltonian. To discuss SPTs in the mixed state context, we need to find an appropriate mixed state generalization. While there are multiple notions that have appeared in literature[18, 51], in our case, if $\rho$ is the density matrix of a mixed state defined on a Hilbert space $\mathcal{H}_{A}$, we say it is SRE if there exists a SRE pure state |$\ket{\psi_{\text{AB}}}\in\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ such that $\rho=\text{Tr}_{B}(\ket{\psi_{\text{AB}}}\bra{\psi_{\text{AB}}})$. In other words, $\rho$ is SRE if it has an SRE purification.
@@
where the second step comes from the clustering of correlations in SRE states and in the last step we used the fact that $\langle q_{-a}^{R}\rangle_{\psi_{\text{AB}}}=\langle\mathcal{I}_{0}q_{-a}^{R}\mathcal{I}_{0}\rangle_{\psi_{\text{AB}}}=e^{i\Theta_{a}}\langle q_{a}^{L}\rangle_{\psi_{\text{AB}}}$. Consequently, the sign of the parity string for odd-length regions is a topological invariant characterizing the state. Note that, in principle, for every pair of inversion related sites there is a topological invariant $\Theta_{a}$. However, in the presence of a translation symmetry (e.g. the two-site translation that arises if site-centered inversion is a symmetry around any site), only two invariant remain [61].
Approximate charge conjugation symmetry and Haldane insulator extended data

### v2

@@ III Coupling an even number of chains removes the topological transition
which results in a gapped path between the stacks of two CSPT-$1$ and CSPT-$2$ phases (Fig. 3a, top). Here, $a_{x,\mathsf{T}/\mathsf{B}}^{(\dagger)}$ labels the bosonic annihilation (creation) operator in the top/bottom layer. Crucially, the above term preserves the protecting $\mathbb{Z}_{2}$ parity symmetry generated by Eq. (2), as well as site-centered inversion.
@@
We simulate the pairing term using a particle-hole mapping, as illustrated in Fig. 3a and detailed in the SM. Starting deep in the CSPT-1 states of both chains with negligible $t_{\perp}$, we ramp across the critical point of a CSPT chain, $\mu_{S}=0$, into a regime where the ground state of each decoupled chain is the CSPT-2 state before imaging. The numerical contour plot in Fig. 3b shows that, with increasing $t_{\perp}$ between two chains, the many-body energy gap between the ground and the first excited states increases at $\mu_{S}=0$, removing the quantum criticality. The ramps in $\mu_{S}$ and $t_{\perp}$ follow the differently shaded red lines at a constant rate for the $\mu_{S}$ ramp.
@@
At low interchain tunneling $t_{\perp}\rightarrow 0$, the ramp cannot be adiabatic as we cross the vanishing energy gap at the quantum critical point. The system heats up, resulting in low fidelity of the final state with respect to the CSPT-2 ground state. This is revealed by measurement of the CSPT-2 order parameter $p_{2}$ (Fig. 3e), which is much smaller than the ground state expectation for small $t_{\perp}$. Similarly, the CSPT-1 order parameter does not vanish in this regime. In contrast, at larger values of $t_{\perp}$, even though the total path traveled in the phase diagram is longer—because we have to change $t_{\perp}$ while simultaneously changing $\mu_{S}$ for the same total sweep time—we observe a more faithful preparation of CSPT-2 state as seen by a larger value of its order parameter.
Figure 4: Mixed-state CSPT from local disorder. a, To demonstrate the importance of site-centered symmetry for protecting the CSPTs, we intentionally add different programmable chemical potential disorder, changing the original light gray lattice to the purple lattice. In the lower panel, the order parameters of a single chemical potential disorder realization (bootstrapping standard error) no longer show clear and distinct phases. The left side lower panel shows the order profile in space for different staggered potentials, while the right side lower panel integrates over the space to plot string order parameters. b, The histograms of the distribution of the $p_{1}$ and $p_{2}$ are plotted for the different disorder realizations. The disorder causes strong deformation in the order parameters for each disorder realization. c, Since the CSPTs are protected by average symmetry, once we average over the different disorder realizations, we obtain a clean signal of two distinct mixed-state phases (standard error of the mean) as seen from the order parameters in the top panel. The order parameters do not reach exactly zero due to finite imaging fidelity, finite size effects, and the degree of approximation of the particle-hole symmetry (SM), but still can distinguish between the two phases. The topological invariant is precisely the sign of the averaged parity string of odd length, seen to be positive for large $\mu_{S}$ and negative for small $\mu_{S}$ in the bottom panel.
The larger energy gap allows us to ramp across the phase diagram with less heating, which is also reflected in the final temperature of the prepared CSPT-2 states (Fig. 3c). The energy of the final state can be well approximated by the energy of the classical state because, at the readout stage, the intrachain tunneling ($t$) is much smaller than other energy scales ($\mu_{S}$, $U$, and $V$). Based on each snapshot, we compute the energy of the final state and obtain a histogram of the state’s distribution. We then fit a temperature and observe the shift towards lower temperature, as $t_{\perp}$ is increased.
@@ IV Mixed State Order from Average Symmetry
Thus far, we have demonstrated the presence of two distinct CSPTs in our experiment and exhibited a hallmark feature of the transition between them, i.e., the disappearance of the transition upon stacking. Crucially, our results depended on the presence of both protecting symmetries of the CSPT—the crystalline site-centered inversion symmetry and the boson parity symmetry—as seen by the fact that the invariant was computed from a parity string of a site-centered inversion symmetric region (namely, one of odd length). To further substantiate our claim that crystalline site-centered inversion symmetry is essential, we explicitly break this symmetry by implementing a programmable chemical potential disorder pattern shown at the top of Fig. 4a, whose magnitude is comparable to the largest term in the Hamiltonian (peak-to-peak $\approx U$, see SM).
In our experiment, we adiabatically prepare the ground state of the disordered staggered dipolar Bose-Hubbard model $t\ll U,\mu_{S}$ by ramping down the tunneling from a superfluid while ramping up the chemical potential disorder pattern. For any given realization of the disorder, we repeat the experiment thousands of times (SM) to sample the wavefunction. This yields the parity string order parameters $p_{1}$ and $p_{2}$ shown in Fig. 4a bottom, which show no signature of a transition with varying $\mu_{S}$, indicating that the transition is removed when inversion symmetry is broken.
However, let us now consider averaging over disorder realizations (SM). We find that the transition is highly sensitive to whether or not we use the knowledge of the disorder. Indeed, we found that when the disorder pattern is known, the transition disappears. However, if this knowledge is erased through averaging—and hence the system forms a mixed state ensemble $\rho$ of the different disordered ground states (SM)—the transition re-emerges [4].
@@
Although for each realization there is no sharp distinction between regimes $\mu_{S}\ll U$ and $\mu_{S}\gg U$ (as seen by the spread in parity values present in the histograms in Fig. 4b), upon averaging over disorder, a sharp distinction is restored (Fig. 4c). This is because the crystalline SPTs we study are also average SPTs, i.e., topological phases of the mixed states that arise due to the randomness of the disorder [41, 42, 43, 44, 4, 45, 46, 47, 48, 49, 50, 51, 52, 53]. Such average SPTs can be shown in theory to be protected by the $\mathbb{Z}_{2}$ parity symmetry along with an average site-centered inversion symmetry of the mixed state corresponding to ^$\hat{\mathcal{I}}\rho\hat{\mathcal{I}}^{\dagger}=\rho$, where $\hat{\mathcal{I}}$ is the unitary operator implementing the action of inversion. Our observations suggest that symmetry protection in realistic quantum systems can persist even in the presence of symmetry-breaking disorder, opening up the exploration of mixed-state phases.
V The Haldane-to-Mott Insulator Transition
We conclude by presenting evidence for the Haldane SPT phase transition in our system and elucidating its connection to the CSPT transition explored in this work. The dipolar Bose–Hubbard model is predicted to host a Mott insulator (MI), separated from the Haldane insulator (HI) [26, 54, 55, 1] by a quantum phase transition when crystalline bond-centered inversion and parity symmetries are present at $\mu_{S}=0$, as shown in Fig. 5a. The HI is a CSPT phase characterized by protected quantum entanglement [10].
@@ V The Haldane-to-Mott Insulator Transition
Naturally, CSPT-1 and MI correspond to the same boson pattern and hence the same phase of matter. More intriguingly, the CSPT-2 phase we encountered for $\mu_{S}\neq 0$ adiabatically connects to HI as we take $\mu_{S}\to 0$ for suitable $U$ and $V$ [3]. The HI can be viewed as a state where bosons are delocalized across lattice bonds, making it symmetric under parity as well as site- and bond-centered inversion symmetries (Fig. 5a; see SM for the model wave function). If we turn on the staggering $\mu_{S}\neq 0$, these delocalized bosons gradually move together to form the CSPT-2 state as schematically shown in Fig. 1a. Moreover, the direct topological phase transition between CSPT-1 and CSPT-2 (for $\mu_{S}\neq 0$) is in the same universality class as the transition between MI and HI (for $\mu_{S}=0$) [38]. We can thus explore the same topological phase transition in two very different energetic regimes.
@@
This is similar to the parity string operator used for CSPT-1 and -2, but with the addition of endpoint dressing $\delta n=n-1$. This modification ensures that Eq. (5) remains nonzero, and its positive value is consistent with CSPT-2 phenomenology (SM). The hallmark signature of the Haldane insulator is then a contrast between the long-range order of $\mathcal{S}_{x_{0}}(r)$ along with the parity string operator $\mathcal{P}_{x_{0}}(r)$ decaying to zero at long distances.
To study the transition between the HI and MI, we turn off the staggered potential and adiabatically load our atoms into the lattice while setting magnetic fields to probe different on-site interactions $U$ [58]. Since snapshots where $\delta n=0$ limit the magnitude of Haldane string order parameter $\mathcal{S}_{x_{0}}(r)$ and it oscillates as a function or $r$, we normalize it as $\widetilde{\mathcal{S}}=(-1)^{r}\eta\mathcal{S}$ with $\eta=1/\langle|\delta n_{i}||\delta n_{i+r}|\rangle$ [17] computed for the state with the maximum string order parameter in the probed parameter space ($V/t\approx 3$). The result is shown in Fig. 5b and reveals two regimes of the phase diagram, one with Haldane-like correlations and the other with Mott-like correlations. In our finite-size system (10 sites), the order parameters are not expected to exhibit sharp features due to large correlation length (SM). Nevertheless, the order parameters we observe are consistent with the simulated results for our finite-size system.
Figure 5: Haldane-to-Mott insulator transition. A well-known SPT phase in the dipolar Bose-Hubbard model is the Haldane Insulator (HI), which can be connected to the CSPT-2 phase. a, Once the staggered field is turned off, the CSPT-1 and -2 phases become Mott and Haldane insulators in the presence of off-site interactions ($V$). A schematic wavefunction for the HI described in Eq. (17) is shown in the inset. b, We tune the on-site interaction $U$ without the staggered field $\mu_{S}$ at fixed $V/t\approx 3$. The order parameters show good agreement with simulated results. The vertical error bars are obtained from bootstrapping standard error; the horizontal error bars are estimated standard error.
@@ VI Conclusion and outlook
Our study reveals that the topological criticality at the transition between the Haldane insulator (HI) and the Mott insulator (MI) can be accessed when probed between crystalline symmetry-protected topological (CSPT) phases. This underscores the power of tunable quantum simulators, which grant access to a broad parameter space and enable the study of quantum criticality in otherwise inaccessible regimes. While local observables fail to distinguish these phases away from the atomic limit, we identified nonlocal string order parameters for the HI-MI and CSPT transitions. Both transitions are predicted to exhibit stacking behavior, which defines the group structure of SPTs and affirms their invertible nature. We demonstrated this property through dynamic ramping across the CSPT critical point. In the CSPT setting, we also observed the emergence of mixed state order, which generalizes topological invariants into mixed states.
Our uncovering of quantum phase transitions between 1D phases admitting atomic limits raises the possibility of analogous phenomena in higher dimensions—an open direction for both theoretical and experimental exploration. These transitions may help chart adiabatic pathways between topologically trivial and non-trivial phases, motivating novel approaches to the discovery of phases of matter [59] starting from experimentally accessible product states such as paramagnets [60, 61]. In parallel, our demonstration of mixed-state order raises compelling questions about its stability, the nature of intervening phases, and how mixed-state perspectives might refine our understanding of real-world quantum systems. Concretely, this same approach could be used to realize topological phases with average crystalline symmetries which can only exist in the presence of disorder [45, 62]. Moreover, it would be interesting to compare the state preparation fidelity of SPT phases via explicit symmetry breaking versus the stacking property demonstrated in our work. The latter might be especially useful for preparing two-dimensional SPT phases protected by, e.g., particle parity symmetry [63], which would be difficult to break explicitly. Our results emphasize that even seemingly simple models can harbor nontrivial topology, reinforcing the idea that topological tools are not just abstract concepts but practical frameworks for identifying and organizing quantum phases.
@@
Figure 6: Chemical potential. a, Our microscope allows three types of potential projections: we project superlattices using the blue accordion lattice, project arbitrary potentials using DMD, and project relatively sharp repulsive walls with blue-detuned elliptical beams. b, A schematic chemical potential landscape along the 1D chain for a 10-site system. With DMD to compensate for harmonic confinement and a pair of repulsive walls shown in chestnut color, the central 10 sites (shaded in purple) give us a well-defined system. The rising wall marked in purple dots follows the chemical potential listed in the text. c, To increase the data acquisition rate, we parallel simulate multiple copies of chains. In the top subfigure, we raise the lattice deep to reduce interchain tunneling and tilt the magnetic field to minimize interchain dipole-dipole interactions. In the bottom subfigure, we use an additional superlattice (light blue) to create isolated pairs of chains.
@@
Fig System size (sites) Atom number required Atom number DS Photon count DS Spin-half mapping DS Overall DS Smallest sample size Average sample size Total sample size
@@
Mapping for coupled-chain simulations
To perform a quantum simulation of the $H_{\perp}$ coupling in main text Eq. (4) in an atom-number-conserving system, we map the unity-filling CSPTs to the half-filling ones and perform a particle-hole transformation on the bottom layer, treating the physical bosons present in the bottom layer as holes (main text Fig. 3a bottom). In this representation, the pairing term ($\Delta_{\perp}$) becomes an inter-layer hopping term ($t_{\perp}$), which can be implemented experimentally. The staggered potential is created by a diagonal superlattice, resulting in a checkerboard pattern.
@@
In the main text, we presented that decoupled chains have two topologically distinct CSPTs, but two tunnel-coupled chains trivialize the topology. Using Exact Diagonalization (ED) (QuSpin [79]) with three by eight sites and four by six sites, we further numerically simulate higher numbers of coupled chains to demonstrate the $\mathbb{Z}_{2}$ properties as shown in Fig. 10.
@@
In this section, we provide some additional details for the theoretical analysis behind our experimental results. While the section below leaves the present work self-contained, an extensive and pedagogical theoretical accompaniment is provided in the forthcoming [38] manuscript.
@@
We start by making a brief remark on the nature of crystalline symmetries for finite systems. Strictly speaking, the system only has a site-centered inversion symmetry for odd-length systems. Nevertheless, for the bulk phase of matter properties, this distinction between even and odd system sizes should not play a role.
CSPT Topological invariant
To get an intuitive understanding of the topological invariant that distinguishes CSPT-1 and CSPT-2, let us first work with the product states that appear in the atomic limit. One way to distinguish these states is to recognize that in any large inversion-symmetric region, the parities of bosons in CSPT-1 and 2 are $-1$ (odd) and $+1$ (even) respectively (main text Fig. 2a for CSPT-1, top). As quantum fluctations are introduced and we move away from the atomic limit, the bosons in these states delocalize into dressed quasiparticles (main text Fig. 2a, bottom) and the expectation value of the parity in this region is no longer exactly $+1$ or $-1$. Nevertheless, the sign of the expectation value of the parity in these regions is a discrete invariant that can be used to distinguish the two. As long as the gap stays open and both parity and inversion symmetry are preserved, any change in the parity expectation within the region occurs symmetrically from its edges. As a result, the sign of the bulk parity remains unchanged (main text Fig. 2a, bottom). The technical tools required to derive this invariant are in Ref. [3] and follows closely the derivation of the mixed state invariant provided later in the SM.
@@
The Landau paradigm characterizes phase transitions by local order parameters that signal spontaneous symmetry breaking [80, 81]. In sharp contrast, the transition we study lies outside this framework: it separates two symmetry-protected topological (SPT) phases that are indistinguishable by any local observable. While local observables such as the average site occupation respond to the staggered chemical potential, they cannot serve as order parameters for distinguishing the two crystalline SPTs. This is because the Hamiltonian already breaks translational symmetry explicitly, so density modulations are always present—even in trivial phases—and they vary smoothly across the transition. To serve as an order parameter, the observable has to be zero outside the phase and non-zero inside the phase. In contrast, the topological invariant distinguishing the phases (e.g., the sign of the inversion-symmetric parity string) changes singularly and is protected by symmetry and a finite gap [3, 10].
@@
The experimental realization of mixed state order can be formalized as follows. We draw disorder realizations $\mathcal{D}$ from a distribution $p_{\mathcal{D}}$ such that disorder patterns related by site-centered inversion are equally likely. Then the ground state for each disorder realization |$\ket{\psi_{\mathcal{D}}}$ is probed by the following
Crucially, probing the above is equivalent to experimentally probing the expected value of the parity string operator in the mixed state $\rho=\sum_{\mathcal{D}}p_{\mathcal{D}}\ket{\psi_{\mathcal{D}}}\bra{\psi_{\mathcal{D}}}$ | with $\mathbb{E}[\mathcal{P}_{x_{0}}(r)]=\text{Tr}\left(\rho\,\hat{\mathcal{P}}_{x_{0}}(r)\right)$.
Now, we prove that the crystalline SPT phases we study persist provided that the site-centered inversion symmetry of the chain is preserved on average. Before delving into the proof, let us review some of the foundational concepts in the many-body physics of mixed states that will be necessary for our analysis.
@@
A central property of pure state SPTs is that they are short-range entangled (SRE)—i.e. they can be prepared from a product state by finite time evolution with a local Hamiltonian. To discuss SPTs in the mixed state context, we need to find an appropriate mixed state generalization. While multiple notions have appeared in literature[4, 83], in our case, if $\rho$ is the density matrix of a mixed state defined on a Hilbert space $\mathcal{H}_{A}$, we say it is SRE if there exists a SRE pure state |$\ket{\psi_{\text{AB}}}\in\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ such that $\rho=\text{Tr}_{B}(\ket{\psi_{\text{AB}}}\bra{\psi_{\text{AB}}})$. In other words, $\rho$ is SRE if it has an SRE purification.
@@
where the second step comes from the clustering of correlations in SRE states, and in the last step we used the fact that $\langle q_{-a}^{R}\rangle_{\psi_{\text{AB}}}=\langle\mathcal{I}_{0}q_{-a}^{R}\mathcal{I}_{0}\rangle_{\psi_{\text{AB}}}=e^{i\Theta_{a}}\langle q_{a}^{L}\rangle_{\psi_{\text{AB}}}$. Consequently, the sign of the parity string for odd-length regions is a topological invariant characterizing the state. Note that, in principle, for every pair of inversion-related sites, there is a topological invariant $\Theta_{a}$. However, in the presence of a translation symmetry (e.g., the two-site translation that arises if site-centered inversion is a symmetry around any site), only two invariants remain [38].
Haldane to Mott insulator phase transition
In this section of the SM, we provide some additional intuition for the connection betwen the Haldane to Mott insulator transition and the transition between CSPT-1 and CSPT-2.